# FAST-Splat: Fast, Ambiguity-Free Semantics Transfer in Gaussian Splatting

### v1

@@ 5 Experiments
We benchmark FAST-Splat against the primary, existing semantic Gaussian Splatting methods LangSplat [26] and Feature-3DGS [49]. Unfortunately, FMGS [51], another primary method, does not provide a publicly-available code implementation, preventing us from benchmarking against this method. We utilize the original code implementation provided by the authors of these baseline methods. In addition, we examine the capability of FAST-Splat in disambiguating the semantic identities of objects, e.g., in scenarios with ambiguous user prompts and scenarios with multiple related objects. We examine each method across five scenes: Kitchen; Cleaning; Meal; Workshop; and Plants. We abbreviate the name of each scene using the initial letter. We utilize YOLO [40] as the closed-set object detector; we provide additional implementation details and further discussion in the Supplementary Material.
Semantic Gaussian Splatting. We examine the fidelity of the $3$D scene reconstruction achieved by FAST-Splat in comparison to prior work LangSplat [26] and Feature-3DGS [49], beginning with a discussion of the computation time required for data pre-processing as well as the training time for each method, reported in Table 1. FAST-Splat requires at least a $13$x faster data pre-processing time compared to the best-competing method Feature-3DGS, and is about $78$x faster when compared to LangSplat. Similarly, FAST-Splat achieves about $5$x (and up to $8$x) faster training times compared to Feature-3DGS, and is about $4$x to $6$x faster to train compared to LangSplat.
Table 1: The data pre-processing time (prep.) and training time (train) for the semantic Gaussian Splatting methods across five scenes. FAST-Splat outperforms all existing methods.
Scene Method Prep. (min) $\downarrow$ Train (min) $\downarrow$ K. LangSplat [26] 128.49 58.54 Feature-3DGS [49] 15.09 73.15 FAST-Splat [ours] 1.12 12.08 C. LangSplat [26] 85.53 73.32 Feature-3DGS [49] 16.49 85.10 FAST-Splat [ours] 1.01 12.44 M. LangSplat [26] 82.67 68.22 Feature-3DGS [49] 16.36 68.01 FAST-Splat [ours] 1.06 11.71 W. LangSplat [26] 103.20 66.32 Feature-3DGS [49] 15.17 65.27 FAST-Splat [ours] 0.81 12.18 P. LangSplat [26] 136.99 62.61 Feature-3DGS [49] 16.05 110.86 FAST-Splat [ours] 1.23 13.79
We utilize standard metrics for comparing $3$D reconstruction methods, such as the peak-signal-to-noise ratio (PSNR) and Structural Similarity Index Measure (SSIM) for evaluating the reconstruction quality; the training time; the memory usage; and the rendering frame-rate. In Table 2, we report the performance of each method on these metrics. FAST-Splat outperforms all other methods on the rendering frame-rate and memory-usage metrics, achieving the highest rendering speeds while requiring the least amount of GPU memory. In four of the five scenes, FAST-Splat achieves the highest SSIM score, outperforming the next best-competing method LangSplat. LangSplat achieves the highest PSNR score in four of the five scenes, followed by FAST-Splat. FAST-Splat achieves the highest PSNR score in the Plants scene. Further, FAST-Splat reduces the memory usage in semantic Gaussian Splatting by a factor of more than $3$x compared to the best-competing method LangSplat and a factor of more than $6$x compared to Feature-3DGS. In essence, FAST-Splat achieves a higher-fidelity scene reconstruction with a lower memory usage.
Lastly, we present the rendering speed of each method. FAST-Splat attains $18$x to $75$x faster rendering speeds compared to the best-competing method LangSplat. Similarly, FAST-Splat achieves about $550$x faster-rendering frame-rates compared to Feature-3DGS, highlighting the superior performance of FAST-Splat. We show rendered images from each of the scenes with the different methods in Figure 2, juxtaposed with the ground-truth RGB image. We render novel views in FAST-Splat.
Table 2: The rendering frame-rate (Speed), image quality, and memory usage (Memory) achieved by the semantic Gaussian Splatting methods across the five scenes.
Scene Method Speed $\uparrow$ (F/s) SSIM $\uparrow$ PSNR $\uparrow$ Memory $\downarrow$ (GB) K. LangSplat [26] 2.25 0.966 35.26 8.05 Feature-3DGS [49] 0.23 0.931 30.20 7.72 FAST-Splat [ours] 115.23 0.971 34.24 1.02 C. LangSplat [26] 2.25 0.918 32.51 9.54 Feature-3DGS [49] 0.22 0.871 28.70 7.72 FAST-Splat [ours] 114.37 0.929 30.35 1.00 M. LangSplat [26] 2.24 0.933 33.37 9.54 Feature-3DGS [49] 0.22 0.845 27.83 7.86 FAST-Splat [ours] 127.41 0.940 31.25 0.98 W. LangSplat [26] 1.87 0.921 31.41 10.08 Feature-3DGS [49] 0.22 0.357 8.11 7.72 FAST-Splat [ours] 141.48 0.917 29.22 1.04 P. LangSplat [26] 6.84 0.915 28.79 7.09 Feature-3DGS [49] 0.21 0.773 25.33 7.72 FAST-Splat [ours] 128.77 0.942 31.12 1.17
@@
Figure 2: RGB images rendered using each semantic Gaussian Splatting method. FAST-Splat achieves $18$x to $75$x faster rendering speeds with at least $3$x lower memory usage, while achieving competitive (and in some cases, better) reconstruction quality.
Semantic Object Localization. Now, we evaluate each method on semantic object segmentation from open-vocabulary queries, using the mean intersection-over-union (mIoU) and accuracy metrics. In Table 3, we report the performance of each method on these metrics. FAST-Splat achieves the best performance in two scenes: the Kitchen scene and the Workshop scene. In the Plants scene, Feature-3DGS achieves the highest mIoU score, while LangSplat achieves the highest accuracy. Further, in the Cleaning scene, Feature-3DGS achieves the highest mIoU score, while FAST-Splat achieves the highest accuracy. In all scenes where FAST-Splat does not achieve the best score, FAST-Splat achieves the next-best score, remaining competitive with the best-performing method, which underscores the overall effectiveness of our proposed method. In Figure 3, we show segmentation masks for each of the methods across the five scenes.
Notably, all methods perform poorly in the Workshop scene. We believe this observation may be due to the Workshop scene being mostly out-of-distribution, considering that the training-data distribution of most vision models and vision-language models do not contain industrial objects, such as those found in workshops. Recall that we utilize the YOLO model for closed-set object detection. The YOLO model is trained using the object categories found in the COCO dataset [19], which does not contain common objects found in a workshop, such as a power drill and screwdriver, which is present in the Workshop scene. Moreover, other open-vocabulary object detection methods also struggle to identify industrial objects, e.g., GroundingDINO [22].
Figure 3: Semantic segmentation using each semantic Gaussian Splatting method. FAST-Splat achieves competitive performance compared to existing semantic Gaussian Splatting method, achieving the highest mIoU and accuracy scores in some scenes while being the next best-performing method in the remaining scenes.
Table 3: The mean intersection-over-union (mIoU) and accuracy of the segmentation masks generated by semantic Gaussian Splatting methods across the five scenes.
Scene Method mIoU $\uparrow$ Accuracy $\uparrow$ K. LangSplat [26] 0.525 0.920 Feature-3DGS [49] 0.550 0.920 FAST-Splat [ours] 0.709 0.925 C. LangSplat [26] 0.552 0.850 Feature-3DGS [49] 0.570 0.751 FAST-Splat [ours] 0.517 0.901 M. LangSplat [26] 0.409 0.714 Feature-3DGS [49] 0.571 0.913 FAST-Splat [ours] 0.525 0.904 W. LangSplat [26] 0.054 0.350 Feature-3DGS [49] 0.061 0.327 FAST-Splat [ours] 0.127 0.908 P. LangSplat [26] 0.562 0.900 Feature-3DGS [49] 0.641 0.872 FAST-Splat [ours] 0.570 0.861
@@
Figure 4: FAST-Splat resolves language ambiguity in user-provided natural-language queries in semantic object localization. FAST-Splat identifies the specific semantic class of each relevant object, e.g., a coffee machine and a kettle, when prompted with an ambiguous query, e.g., “coffee” and “cooking pot,” respectively. Likewise, FAST-Splat disambiguates between a “cup” and a vase and between a “fruit” and a pottedplant in the garden-like scene.
@@
Semantic Disambiguation. FAST-Splat extends the capabilities of semantic Gaussian Splatting methods beyond semantic object localization to the domain of semantic disambiguation. As noted in the preceding discussion, existing semantic Gaussian Splatting methods cannot resolve semantic ambiguity either arising from the user-provided natural-language query or inherently from the scene composition. In contrast, when provided with an ambiguous natural-language query, FAST-Splat disambiguates the semantic object localization task, by identifying the specific semantic label of related objects in the scene.
@@
We provide some examples in Figure 4. For example, when a user queries for “coffee” in the Kitchen scene, FAST-Splat provides the user with the semantic identity of the related object coffee machine, along with the semantic segmentation mask and location of the coffee machine, given a threshold for the segmentation task. Likewise, when queried for a “cooking pot,” FAST-Splat identifies the semantic class: kettle, and the segmentation mask and location for the kettle. Notably, in the Plants scene, FAST-Splat resolves the ambiguity between a cup and a vase. When queried with the prompt “cup,” FAST-Splat informs the user that, more precisely, a vase exists in the scene. Similarly, when asked to localize a “fruit,” FAST-Splat clarifies that the scene contains a pottedplant and provides the semantic segmentation mask and location.
@@
Scene Editing. Further, like prior semantic Gaussian Splatting methods, FAST-Splat enables scene-editing of Gaussian Splatting scenes. With FAST-Splat, a user can modify the visual properties of objects in the scene, and in addition, can insert, delete, or move objects in the scene. For space considerations, we omit a detailed discussion of this capability, given that it has been discussed in prior work. In Figure 5, we provide an example where the visual properties of the coffee machine is modified.
Figure 5: Like existing semantic Gaussian Splatting methods, FAST-Splat enables editing of Gaussian Splatting scenes. Here, we show color-editing of a coffee machine in a kitchen.
@@ 6 Conclusion
We introduce FAST-Splat, a fast semantic Gaussian Splatting method that resolves ambiguity in semantic object localization. FAST-Splat retains the explicit form of Gaussian Splatting and avoids neural architectures to achieve fast training and rendering speeds compared to existing methods. Further, FAST-Splat builds upon closed-set semantic distillation to enable open-vocabulary semantic Gaussian Splatting. We leverage this formulation to endow FAST-Splat with the capability to resolve semantic ambiguity arising from the natural-language query or the scene composition, in the semantic object localization task, enabling FAST-Splat to address one of the fundamental limitations of existing semantic Gaussian Splatting methods.
@@ 7 Limitations and Future Work
As with existing semantic Gaussian Splatting methods, the performance of FAST-Splat is highly impacted by the detection performance of the closed-set object detector. Currently, the size of the object categories detected by open-source closed-set object detectors is quite limited. For example, YOLO [40] can only detect $80$ objects, which is quite small, considering the diversity of objects in the real-world. This limitation is not imposed by the underlying model architecture. In fact, YOLO can be trained on a much greater diversity of objects. However, the model was trained using the COCO dataset, introducing this limitation. In comparison, existing vision-language models are trained on more diverse datasets, enabling these models to detect a much broader range of objects. Future work will seek to utilize closed-set object detectors trained with a much larger set of object categories.
Our work utilizes an open-vocabulary object detector, e.g., GroundingDINO [22], in addition to a closed-set object detector, which enables FAST-Splat to mitigate the limitations of the closed-set object detector. Essentially, we can always augment the set of detectable objects through natural-language when querying the open-vocabulary object detector. In fact, we are currently exploring using image-tagging models, e.g., RAM [47], or multi-modal vision-language models, e.g., LLaVa [20] to detect objects for language grounding. Future work will seek to explore this research direction.
@@ Appendix A Implementation Details
We collected the training data by taking videos of the scene using a smartphone camera. We utilized structure-from-motion, e.g., COLMAP [33], for computing the pose of the camera for each frame used in training the Gaussian Splatting model. FAST-Splat is trained on an NVIDIA GeForce RTX 3090 GPU with 24GB VRAM for $30000$ iterations, using the hyperparameters specified in [38] in its Splafacto model. We train the baseline methods LangSplat and Feature-3DGS using the original code implementation provided by their respective authors on an NVIDIA GeForce RTX 3090 Ti GPU with 24GB VRAM, giving LangSplat and Feature-3DGS an advantage on hardware, since the 3090 Ti has $10752$ CUDA cores compared to the $10496$ CUDA cores on the 3090, among other improved hardware specifications. To train LangSplat, we follow the procedure provided by its authors, which involves pre-training the Gaussian Splatting Scene using the code implementation provided in [11], prior to training the semantic components. We do not include the structure-from-motion processing times in any of the results reported in the experiments, given that this procedure is required universally by all methods.
Appendix B Semantic Object Localization

### v2

@@ 5 Experiments
We examine the semantic segmentation and disambiguation capabilities of FAST-Splat, e.g., in scenarios with ambiguous user prompts and scenarios with multiple related objects. We summarize the results here and provide additional implementation details and results in the Supplementary Material.
Baselines. We evaluate FAST-Splat against existing semantic GSplat methods: LangSplat [26] and Feature-3DGS (F-3DGS) [50]. We do not compare our work with FMGS [52], which does not have a publicly-available code implementation. We utilize the original code implementation provided by the authors of these baseline methods.
Datasets and Metrics. We evaluate each method across the benchmark datasets: 3D-OVS [20], LERF [13], and MipNeRF360 [2]. Further, we examine each method in uncurated real-world scenes, e.g., in typical household settings. In each dataset, we report the mean intersection-over-union (mIoU), localization accuracy, and training time [mins].
5.1 Comparison: Benchmark Datasets.
Figure 2: Semantic segmentation on the Bed and Covered-Desk scenes in 3D-OVS. Overall, FAST-Splat outperforms the baselines.
We evaluate all methods on the 3D-OVS dataset, composed of a set of long-tail objects in different backgrounds and poses, and the LERF dataset. Table 1 presents the mIoU scores achieved by each method in the 3D-OVS dataset, with FAST-Splat achieving the highest mIoU in each scene. Specifically, FAST-Splat outperforms LangSplat and F-3DGS by about $8.3\%$ and $40\%$, respectively, in mIoU. Similarly, from Table 2, FAST-Splat achieves the highest localization accuracy scores, outperforming LangSplat and F-3DGS by about $1.2\%$ and $7.4\%$, respectively.
Moreover, the novel architecture of FAST-Splat enables significantly faster training, highlighted in Table 3, where FAST-Splat achieves $5.6$x and $7.7$x speedups in training time compared to LangSplat and F-3DGS, respectively. LangSplat requires two independent training phases for each scene (one for the underlying Gaussian representation and another for the distillation procedure), resulting in significantly longer training times. In addition, LangSplat utilizes costly data pre-processing procedures, making multiple queries to the CLIP and SAM models for semantics extraction, further degrading its computational performance. Similarly, F-3DGS suffers from these challenges. In contrast, FAST-Splat overcomes these challenges with its single-phase training procedure (see Figure 1), resulting in much faster training times. These results underscore the superior performance of FAST-Splat compared to existing methods. In Figure 2, we show the semantic segmentation masks generated by each method. Unlike F-3DGS, which fails to segment many of the objects in the scenes, FAST-Splat and LangSplat provide quite accurate segmentation masks. However, on average, FAST-Splat achieves higher accuracy. We present the results on the LERF dataset in Appendix B.
Table 1: mIoU Scores in the 3D-OVS Dataset.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] Bed 0.925 0.643 0.933 Covered-Desk 0.779 0.678 0.913 overall 0.852 0.661 0.923
Table 2: Localization accuracy in the 3D-OVS Dataset.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] Bed 0.992 0.961 0.998 Covered-Desk 0.978 0.895 0.995 overall 0.985 0.928 0.997
Table 3: Total training times [mins] in the 3D-OVS Dataset.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] Bed 79.23 107.25 13.96 Covered-Desk 78.36 112.06 13.96
5.2 Ablations.
Next, we ablate the choice of using an open-set detector compared to using only a closed-set detector in FAST-Splat using the MipNerf360 datasets. We use the YOLO model for closed-set object detection. In Tables 5 and 5, we report the mIoU and localization accuracy associated with each approach. From these results, the open-set detector significantly improves the segmentation performance of FAST-Splat beyond that of the closed-set detector. In particular, the open-set detector improves the overall mIoU and localization accuracy scores by about $97\%$ and $25\%$, respectively, compared to the closed-set detector. This performance boost can be explained by the greater expressiveness of the open-set detector, enabling it to identify varied objects, compared to the relatively limited set of objects that can be identified by a closed-set detector. However, using an open-set detector results in slightly longer training times. The closed-set approach requires about $10.30$ mins and $25.73$ mins in the Kitchen and Bicycle scenes, respectively, compared to the open-set approach, which requires about $12.28$ mins and $27.82$ mins in the Kitchen and Bicycle scenes. We provide further ablation results, along with renderings of the segmentation masks, using the 3D-OVS and LERF datasets in Appendix C.
Table 4: IoU Scores in MipNeRF360.
Scene Closed-Set Open-Set Kitchen 0.366 0.869 Bicycle 0.395 0.633 overall 0.381 0.751
Table 5: Localization accuracy in MipNeRF360.
Scene Closed-Set Open-Set Kitchen 0.887 0.976 Bicycle 0.577 0.855 overall 0.732 0.916
@@
Figure 3: RGB images rendered in R-Kitchen and R-Library. FAST-Splat achieves $18$x to $51$x faster rendering speeds with at least $6$x lower memory usage, while achieving competitive or better reconstruction quality.
5.3 Comparison: Uncurated Real-World Datasets.
In addition, we create a dataset of images from uncurated real-world scenes to capture the diversty of objects that would typically occur in real-world scene. We collect data from (i) a home kitchen in its natural, everyday state, comprising the R-Kitchen dataset and (ii) a home library, containing books and plants, comprising the R-Library dataset. We keep the scale of the training images in the uncurated dataset the same as that of MipNerf360, a widely-used benchmark dataset. We describe the data setup in greater detail in the Appendix A. Here, we examine the fidelity of the $3$D scene reconstruction, reporting the structural similarity index measure (SSIM), the peak signal-to-noise ratio (PSNR), the rendering speed in frames per second [FPS], and memory usage [GB] in Table 6. FAST-Splat enables real-time rendering speeds, essentially retaining the performance of the original GSplat work [12], compared to LangSplat and F-3DGS, whose performance is limited by their semantics distillation design. FAST-Splat achieves $51$x and $18.8$x speedup in rendering speed and the highest SSIM scores compared to the best-competing method, in the R-Kitchen and R-Library datasets, respectively. LangSplat achieves the highest PSNR score in the R-Kitchen dataset, while FAST-Splat achieves the highest PSNR score in the R-Library dataset. Further, FAST-Splat reduces the memory usage by about $7.6$x and $6.1$x and the training times by about $6.7$x and $8.4$x, compared to the best competitor in both datasets.
Table 6: The rendering frame-rate (Speed [FPS]), image quality, and memory usage (Memory [GB]) in Uncurated Real-World Datasets.
Method Speed $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ Memory $\downarrow$ R-Kitchen LangSplat [26] 2.25 0.966 35.26 8.05 F-3DGS [50] 0.23 0.931 30.20 7.72 FAST-Splat [ours] 115.23 0.971 34.24 1.02 R-Library LangSplat [26] 6.84 0.915 28.79 7.09 F-3DGS [50] 0.21 0.773 25.33 7.72 FAST-Splat [ours] 128.77 0.942 31.12 1.17
Table 7: Total training times [mins] in Uncurated Real-World Datasets.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] R-Kitchen 187.03 88.24 13.20 R-Library 199.60 126.91 15.02
@@
Figure 4: FAST-Splat resolves language ambiguity in natural-language queries in semantic object localization, identifying the specific semantic class of each object, e.g., a coffee machine and a kettle, when prompted with an ambiguous query, e.g., “coffee” and “cooking pot,” respectively. Likewise, FAST-Splat disambiguates between a “cup” and a vase and between a “fruit” and a pottedplant in R-Library.
@@
Semantic Disambiguation. FAST-Splat extends the capabilities of semantic GSplat methods beyond semantic object localization to the domain of semantic disambiguation. As noted in the preceding discussion, existing semantic GSplat methods cannot resolve semantic ambiguity either arising from the natural-language query or inherently from the scene composition. In contrast, when provided with an ambiguous natural-language query, FAST-Splat disambiguates the semantic object localization task, by identifying the specific semantic label of related objects in the scene.
@@
We provide some examples in Figure 4. For example, when a user queries for “coffee” in the R-Kitchen scene, FAST-Splat provides the user with the semantic identity of the related object, coffee machine, along with its semantic segmentation mask and location, given a threshold for the segmentation task. Likewise, when queried for a “cooking pot,” FAST-Splat identifies the semantic class: kettle, including its segmentation mask and location. Notably, in the R-Library scene, FAST-Splat resolves the ambiguity between a cup and a vase. When queried with the prompt “cup,” FAST-Splat informs the user that, more precisely, a vase exists in the scene. Similarly, when asked to localize a “fruit,” FAST-Splat clarifies that the scene contains a pottedplant and provides the semantic segmentation mask and location.
@@
Scene Editing. FAST-Splat enables scene-editing of GSplat scenes, by leveraging semantics in Gaussian Splatting. With FAST-Splat, a user can modify the visual properties of a scene, and in addition, insert, delete, or move objects in the scene. For space considerations, we omit a detailed discussion of this capability, given that it has been discussed in prior work. In Figure 5, we provide an example where the color of the coffee machine in R-Kitchen is modified.
Figure 5: Color-editing of a coffee machine using FAST-Splat.
@@ 6 Conclusion
We introduce FAST-Splat, a fast semantic GSplat method that resolves ambiguity in semantic object localization. FAST-Splat retains the explicit form of Gaussian Splatting and avoids neural architectures to achieve fast training and rendering speeds compared to existing methods. FAST-Splat jointly trains the geometric, visual, and semantic attributes of a scene without any neural models, resulting in faster training times, rendering rates. FAST-Splat leverages a hash-table of semantic codes to resolve semantic ambiguity arising from the natural-language query or the scene composition, in the semantic object localization task, addressing one of the fundamental limitations of prior semantic GSplat methods.
@@ 7 Limitations and Future Work
FAST-Splat relies on 2D vision-language foundation models, e.g., CLIP, for distilling semantics into GSplat scenes. Hence, the quality of the resulting semantic GSplat depends on the expressiveness of the semantic embeddings computed by the vision-language model. FAST-Splat might fail to recognize objects if the vision-language model fails to correctly identify such objects. Future work will seek to estimate the uncertainty of the semantic embeddings computed by vision-language models to better characterize the limitations of the trained semantic GSplat and ultimately provide safety and performance guarantees for the semantic GSplat.
@@ Appendix A Implementation Details
For the uncurated real-world datasets, we collected the training data by taking videos of the scene using a smartphone camera. We utilized structure-from-motion, e.g., COLMAP [33], for computing the pose of the camera for each frame used in training the GSplat model. FAST-Splat is trained on an NVIDIA GeForce RTX 3090 GPU with 24GB VRAM for $30000$ iterations, using the hyperparameters specified in [39] in its Splafacto model. We train the baseline methods LangSplat and Feature-3DGS using the original code implementation provided by their respective authors on an NVIDIA GeForce RTX 3090 Ti GPU with 24GB VRAM, giving LangSplat and Feature-3DGS an advantage on hardware, since the 3090 Ti has $10752$ CUDA cores compared to the $10496$ CUDA cores on the 3090, among other improved hardware specifications. To train LangSplat, we follow the procedure provided by its authors, which involves pre-training the Gaussian Splatting Scene using the code implementation provided in [12], prior to training the semantic components. We do not include the structure-from-motion processing times in any of the results reported in the experiments, given that this procedure is required universally by all methods.
Appendix B Comparison on the LERF Dataset
We provide additional results on the segmentation performance of each method in the LERF dataset. In Tables 8, 9 and 10, we provide the mIoU and localization accuracy scores achieved by FAST-Splat, LangSplat, and F-3DGS. FAST-Splat achieves the highest overall mIoU score and localization accuracy and achieves the highest localization accuracy in each scene. F-3DGS achieves the highest mIoU score in the Ramen and Waldo_kitchen scenes, while FAST-Splat achieves the highest mIoU scores in the Teatime and Waldo_kitchen scenes. From Table 10, FAST-Splat requires significantly shorter training times compared to the baselines, providing $7.7$x, $6.9$x, and $7.5$x speedups compared to the best-competing methods in the Ramen, Teatime, and Waldo_kitchen scenes, respectively. These results further highlight the superior performance of FAST-Splat compared to existing baselines.
Table 8: mIoU Scores in the LERF Dataset.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] Ramen 0.512 0.600 0.566 Teatime 0.651 0.613 0.706 Waldo_kitchen 0.445 0.533 0.533 overall 0.536 0.582 0.602
Table 9: Localization accuracy in the LERF Dataset.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] Ramen 0.732 0.876 0.935 Teatime 0.881 0.890 0.972 Waldo_kitchen 0.955 0.863 0.971 overall 0.856 0.876 0.959
Table 10: Total training times [mins] in the LERF Dataset.
Scene LangSplat [26] F-3DGS [50] FAST-Splat [Ours] Ramen 141.98 116.49 15.08 Teatime 175.26 114.53 16.70 Waldo_kitchen 189.55 123.77 16.59
Figure 6: Semantic segmentation on the Bed scene in 3D-OVS. Whereas the closed-set approach fails to produce high-quality segmentation masks, the open-set approach achieves high accuracy.
Figure 7: Semantic segmentation on the Covered-Desk in 3D-OVS. The open-set approach achieves higher accuracy, especially for long-tail objects, e.g., the gerbera (a daisy/flower).
Appendix C Ablations.
Further, we discuss ablation results in the 3D-OVS and LERF datasets, where we evaluate the performance of FAST-Splat with only a closed-set detector and the variant with an open-set detector. Tables 11 and 12 provide the mIoU and localization accuracy scores for the closed-set and open-set approaches in the 3D-OVS dataset. In this dataset, the open-set approach improves the mIoU score by about $8.8$x and the localization accuracy by about $1.9$x relative to that of the closed-set approach. In Figures 7 and 7, we show the segmentation masks produced by each approach. The closed-set approach fails to produce meaningful segmentation masks in the Bed scene; in contrast, the open-set approach produces highly-accurate segmentation masks. The closed-set approach performs better in the Covered-Desk scene (relative to the Bed scene), localizing the teddy bear. However, the closed-set approach fails when segmenting long-tail objects such as the gerbera in Figure 7, which the open-set approach successfully localizes.
We observe similar findings in the LERF dataset in Tables 14 and 15, where the open-set approach improves the mIoU and localization accuracy scores by about $2$x and $1.15$x. Figures 8 and 9 shows the segmentation masks generated by each approach. While the closed-set approach fails to produce highly-detailed segmentation masks, e.g., segmenting the egg, noodles, and pork-belly in the bowl, the open-set approach provides higher-fidelity segmentation masks. Likewise, the closed-set approach fails to distinguish the sheep from the teddy bear in Figure 9, unlike the open-set approach. However, in all the datasets, the closed-set approach requires shorter training times compared to the open-set approach.
Lastly, in Figure 10, we provide a few segmentation images generated by the closed-set and open-set approaches in the MipNerf360 datasets. Although the closed-set approach fails to accurately segment the bicycle behind the bench, the open-set approach successfully segments the bicycle. These results underscore the importance of the open-set detector in FAST-Splat.
Table 11: mIoU Scores in the 3D-OVS Datasets.
Scene Closed-Set Open-Set Bed 0.044 0.933 Covered-Desk 0.165 0.913 overall 0.105 0.923
Table 12: Localization accuracy in the 3D-OVS Datasets.
Scene Closed-Set Open-Set Bed 0.165 0.998 Covered-Desk 0.893 0.995 overall 0.529 0.997
Table 13: Total training times [mins] in the 3D-OVS Datasets.
Scene Closed-Set Open-Set Bed 12.6 13.96 Covered-Desk 12.07 13.96
Table 14: mIoU Scores in the LERF Datasets.
Scene Closed-Set Open-Set Ramen 0.135 0.566 Teatime 0.407 0.706 Waldo_kitchen 0.343 0.533 overall 0.295 0.602
Table 15: Localization accuracy in the LERF Datasets.
Scene Closed-Set Open-Set Ramen 0.736 0.935 Teatime 0.916 0.972 Waldo_kitchen 0.858 0.971 overall 0.837 0.959
Table 16: Total training times [mins] in the LERF Datasets.
Scene Closed-Set Open-Set Ramen 13.86 15.08 Teatime 15.01 16.70 Waldo_kitchen 13.99 16.59
Figure 8: Semantic segmentation of the Ramen scene in the LERF dataset. The closed-set approach fails to generate highly-detailed segmentation masks, unlike the open-set approach.
Figure 9: Semantic segmentation of the Teatime scene in the LERF dataset. The closed-set achieves a relatively better segmentation performance compared to other scenes, but fails to distinguish between the sheep and the teddy bear.
Figure 10: Semantic segmentation of the Bicycle scene in the MipNerf360 dataset. The closed-set approach fails to accurately segment the bicycle behind the bench, unlike the open-set approach.
Appendix D Semantic Object Localization